# Roll in the Tanks! Measuring Left-wing Extremism on Reddit at Scale

### v3

@@ Appendix E What are their lexical differences?
Aligning Multidimensional Word Embeddings. Understanding the language usage in polarized communities is challenging due to their distinct vocabularies for identical concepts. E.g., Democrats and Republicans are political adversaries. Thus, a Democrat might say that “Republicans are America’s greatest threat” (and vice versa), making direct comparisons of their language less useful when it comes to polarizing topics. To address this, KhudaBuksh et al. (KhudaBukhsh et al., 2021) formulate the problem as a form of translation between two groups using the same natural language. While not every “translation” is as clear-cut as the previous toy example, this technique has seen refinement (Milbauer et al., 2021) to capture more complex, multidimensional aspects of ideological differences. As such, it emerges as an ideal tool for distinguishing disparities between tankies and other left-wing ideologies.
In a nutshell, we align two multidimensional word embeddings by creating a bilingual lexicon to anchor the alignment. We then use a translation function to project the embeddings from one community to another. If a word from a source community corresponds to a different word in a target community, we consider the words to be “misaligned.” We can further discover conceptual homomorphisms between communities, enabling us to identify equivalent words in the target community for unique words or slang used in the source community. This is accomplished by projecting the words from the source community’s vocabulary that are not in the vocabulary of the target community to target community’s vocabulary. This allows us to get a much deeper understanding of where tankies sit with respect to the rest of the far-left.
@@ E.1.2. Quantative Validation
Appendix F Toxicity Analysis
Perspective API. The Perspective API (Perspective API, 2018) is a widely used (Chong and Kwak, 2022; Balci et al., 2023; Yudhoatmojo et al., 2024) tool for measuring toxicity. Although it has limitations, e.g., there are issues of bias and questions of performance when encountering conversation patterns that it was not trained on, at scale it provides a decent measure for comparison between online communities. The API provides six production models: 1) TOXICITY, 2) SEVERE_TOXICITY, 3) INSULT, 4) IDENTITY_ATTACK, 5) THREAT, and 6) PROFANITY (See (Perspective, 2022) for full details on the models). We consider a threshold of 0.8, defined as “high” for the SEVERE_TOXICITY scores by Hoseini et al. (Hoseini et al., 2023). To have a baseline for the comparisons, we sample 0.5% of the Reddit posts during the dataset’s timeline, which accounts to more than 36 M posts.

### v4

@@ Appendix E What are their lexical differences?
If a word from a source community corresponds to a different word in a target community, we consider the words to be “misaligned.” We can further discover conceptual homomorphisms between communities, enabling us to identify equivalent words in the target community for unique words or slang used in the source community. This is accomplished by projecting the words from the source community’s vocabulary that are not in the vocabulary of the target community to target community’s vocabulary. This allows us to get a much deeper understanding of where tankies sit with respect to the rest of the far-left.
@@ E.1.2. Quantative Validation
Appendix F Toxicity Analysis
Perspective API. The Perspective API (Perspective API, 2018) is a widely used (Aliapoulios et al., 2021; Horta Ribeiro et al., 2021) tool for measuring toxicity. Although it has limitations, e.g., there are issues of bias and questions of performance when encountering conversation patterns that it was not trained on, at scale it provides a decent measure for comparison between online communities. The API provides six production models: 1) TOXICITY, 2) SEVERE_TOXICITY, 3) INSULT, 4) IDENTITY_ATTACK, 5) THREAT, and 6) PROFANITY (See (Perspective, 2022) for full details on the models). We consider a threshold of 0.8, defined as “high” for the SEVERE_TOXICITY scores by Hoseini et al. (Hoseini et al., 2023). To have a baseline for the comparisons, we sample 0.5% of the Reddit posts during the dataset’s timeline, which accounts to more than 36 M posts.